OpenStudy (anonymous):

how do you solve e^(2x) − 4e^x + 3 = 0

OpenStudy (amistre64):

its just a quadratic in disguise

OpenStudy (amistre64):

let u = e^x; x = ln(u) u^2 -4u +3 = 0 once you know u, you know x

OpenStudy (anonymous):

why is x= ln (u) ?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

its a substitution \[(e^x)^2=e^{2x}\] since e^x is just "some unknown value", it is a variable let u = e^x when we know the value of u, say its 5 5 = e^x , we still need to solve for x; ln(u) = x

OpenStudy (amistre64):

im assuming you know what an inverse function is.

OpenStudy (amistre64):

e^x = y, if and only if, x = ln(y) if you do not know this relation yet, then it would take a bit of explaining to prove :)

OpenStudy (anonymous):

thanks I get it :)

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

howdy